OpenStudy (lgbasallote):

You are at a department store with your parents when you saw a $7 shirt. You asked your parents for money to buy the shirt. Your mom gave you $5 and your dad gave you another $5. Therefore, your change is $3.You gave $1 to your mom; $1 to your dad and $1 to yourself. Therefore, you're remaining loan to your mom is $4 and your dad $4 as well. Adding the two $4 will give you $8 and when added to the $1 you have, totals $9. Where did the $1 go? (This is similar to the missing dollar riddle)

OpenStudy (anonymous):

Add properly :P

OpenStudy (lgbasallote):

oh, but whatever do you mean mr. fool? /:)

OpenStudy (experimentx):

tricky question of words... ! you have $8 loan, and you posses $1, if you gave $1 to your parent you will have $7 loan, of which you bought shirt. -$7+shirt = 0, man where you adding, you should subtract $1 from $8

OpenStudy (lgbasallote):

hmmm...i dont quite get your explanation...maybe written in an equation would make me see? yes? :D

OpenStudy (anonymous):

A) Dad and mum gave you 10 rupees B) Your billed amount is Rs. 7 C) You returned you 2 rupees D) You keep 1 rupee Balance is =(A-B)-(C+D) = (10-7) -(2+1) = 0

OpenStudy (experimentx):

you have $1 so you have $8 loan, you CANNOT ADD $1 on this $8 to get $9 ... so you can say that $1 is missing. It's because $8 is loan, i.e. -$8 and you have $1, it's +$1

OpenStudy (lgbasallote):

very good :D